Peggy paid $51.30 for 2 boxes of chocolates, 6 roses, and 3 greeting cards. Joe paid $72.05 for 3 boxes of chocolates, 12 roses, and 2 greeting cards. If 1 greeting card costs the same as 2 roses, how much does a box of chocolates cost
Answer by mananth(16949)

Roses = x
greeting card = 2x
chocolate =y
$51.30
2 boxes of chocolates ---------y each
6 roses,----------------------- x each
3 greeting cards.--------------2x
2y+6x+6x=51.30
2y+12x=51.30....................1
Joe paid $72.05
for 3 boxes of chocolates,------------y each
12 roses,-----------------------------x
2 greeting cards----------------------2x
3y+12x+4x=72.05
3y+16x=72.05....................2
12 x + 2 y = 51.3 .............1
16 x + 3 y = 72.05 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 2
-36 x -6 y = -153.9
32 x 6 y = 144.1
Add the two equations
-4 x = -9.8
/ -4
x = 2.45
plug value of x in (1)
12 x + 2 y = 51.3
29.4 + 2 y = 51.3
2 y = 51.3 -29.4
2 y = 21.9
y = 10.95
chocolates box = $10.95

m.ananth@hotmail.ca